# references.bib

### v2

@@ Results
Recall that the particle filter is capable of combining a mechanistic model of the within-host dynamics, such as the TEIRV, and an observation model, such as the viral load measurements, and will return a (Bayesian) posterior sample of both the parameters of the process and the trajectory of virus and cell populations through time. We assume that the infection begins with a single exposed target cell (in a population of $8\times 10^{7}$ target cells) ke2022daily. This gives us an initial condition for the process: $T(0)=8\times 10^{7}$, $E(0)=1$, $I(0)=0$, $R(0)=0$. We leave the initial viral load, $V(0)$, as a parameter to be fit to the data.
We selected six patient time series ke2022daily, choosing ones that contained a full 14 data points, for both consistency and simplicity. Two of the model parameters were fixed as in previous analysis ke2022daily: $c=10$, $k=4$. See SI Section 4 for full details of the particle filter and JSF configuration.
@@
Figure 4: JSF enables the inference of the viral load through time and the probability that the virus has been cleared from longitudinal data of SARS-CoV-2 viral load measurements. A Model fits to viral load data based on nasal swabs of six patients from ke2022daily using a refractory cell within-host model, using our Jump-Switch-Flow method. Nasal viral load data is shown as solid dots, and the limit of detection is shown as a dotted line. We show the switching threshold, $\Omega=10^{2}$ by a dashed line. The red and purple bands show mean (0%), 25%, 50%, 75% and 95% credible intervals (from dark to light) for the inference and prediction, respectively. We also estimate the probability of viral clearance at a given point in time, based on the previous data to that point, in blue (light). B The posterior distribution of viral $R_{0}$ for each patient, along with the median (white circle), These patient specific estimates are contrasted with the single point estimate from previous analysis ke2022daily.
We estimate the probability of viral clearance (right blue axis), over time, given by the blue lines, where viral clearance is said to occur when all of the virion, eclipsed, and infected cell populations become extinct.
There is clear heterogeneity in the amount of time required to clear the virus (Figure 4A). We estimate that with high certainty patients 423192, 443108 and 445602 clear the virus within 20 days. In contrast, for patients 444332, 444391 and 451152 we infer that they have not cleared the virus within that time frame. Within those who did clear the virus, we observe that different patients require different amounts of time to clear the virus. Specifically, patient 432192 (4A) is inferred to have cleared the virus first. Patients 443108 and 445602 both require an estimated 16 days to obtain viral clearance.
The parameter estimates and priors, are given in SI Section 4. Because of potential identifiability issues with the rate parameters, we instead compare the estimated reproduction number, $R_{0}$, to results from previous analysis ke2022daily. Figure 4B shows the posterior distributions of $R_{0}$, as well as the point estimate of ke2022daily. Since the previous analysis, ke2022daily, reports identical estimates for $\pi$, $\beta$ and $\delta$ across these patients, we only include a single point estimate for their work. Our estimate is consistent with previous results ke2022daily for patients 432192, 445602 and 451152. However, patients 443108, 444332 and 444391 all have higher $R_{0}$ estimates in our results. Our $R_{0}$ estimates are consistent with similar within-host viral infection analyses of other respiratory pathogens (baccam2006kinetics; hernandez2020host; gubbins2024quantifying).
@@ Discussion
We have presented a simple efficient hybrid simulation method for compartmental models: Jump-Switch-Flow (JSF). JSF allows compartments to dynamically change between stochastic and deterministic behaviour. Combining JSF with a simulation based inference method, e.g. a particle filter, enables inference for multi-scale models, in particular when there are atto-type problems where the discrete nature of small populations is important. We demonstrated the properties of JSF in extensive simulation studies and applied it to the analysis of SARS-CoV-2 viral load data.
@@
Simulations shows JSF produces trajectories largely consistent with gold standard exact simulation techniques, e.g. Doob-Gillespie algorithm (see Figs. S4 and S6). However, JSF is much faster for realistic population sizes (see Fig. S7). We demonstrate how our approach can be incorporated into a particle filter (Moss2024), to perform parameter and state estimation and prediction (see Fig. 2). A strength of JSF is the capacity to infer epidemic fade-out, which is important for calibrating intervention measures.
@@
We analysed viral load for SARS-CoV-2 infections using the JSF within a particle filter with a TEIRV, refractory cell model (ke2022daily). In the subset of the data considered, we find consistent parameter values between patients (SI Section 5). In a novel analysis, we estimated the probability of viral clearance through time, finding substantial heterogeneity in the time until viral clearance (see Fig. 4A). This is important as an accurate quantification of when the virus has been cleared is crucial for determining appropriate treatment regimes.
@@
A key advantage of JSF is the ability to combine stochasticity of discrete model for small populations with convenient deterministic models for large populations. The point at which a compartment will transition between these descriptions is specified by the threshold parameters $\Omega_{i}$ (one for each compartment). Setting these parameters to low values speeds up computation, while setting them higher captures more of the stochasticity in the process. As demonstrated in the simulation studies, an appropriate value can be determined through some preliminary simulations. An important consideration in selecting this parameter is ensuring that the absorbing states can still be reached (e.g. compartment extinction), and that stable and steady states can be perturbed by random fluctuations. Furthermore, developing theory to better understand when a CTMC is well approximated by an ODE is at the core to determine how the switching threshold should be chosen. However, this lies outside the scope of this work.
Describing a model via stoichiometric matrices forces systematic consideration of exchange between compartments, encouraging realistic descriptions of physical systems. This formalism can be assisted with Petri nets, which more formally describe such systems (gilbert2006petri) and are an area for future work. Additional extensions include the incorporation of an intermittent Stochastic Differential Equation approximation between the CTMC and ODE regimes, and extension to spatial processes which exhibit both stochastic and deterministic behaviours.
@@ Acknowledgement
We thank Ada Yan for helpful discussions regarding within-host modelling. J.A.F.’s research is supported by the Australian Research Council (DP200100747, FT210100034) and the National Health and Medical Research Council (APP2019093). S.H.’s research is supported by the Australian Research Council (DP200101281).
@@
J.A.F and M.B.F conceptualized research; D.P.J.G, A.E.Z, S.H., R.M, J.A.F and M.B.F designed research; D.P.J.G and A.E.Z performed research; D.P.J.G, A.E.Z and R.M contributed software tools; D.P.J.G and A.E.Z analyzed data; J.A.F and M.B.F supervised research; D.P.J.G and A.E.Z wrote the paper; D.P.J.G, A.E.Z, S.H., R.M, J.A.F and M.B.F reviewed the paper; J.A.F acquired funding.
[…]
[heading=bibintoc]
Supplementary Information: Jump-Switch-Flow: hybrid stochastic-deterministic solutions of compartmental models
@@
S1 Jump-Switch-Flow: mathematical details
In this section, we shall address the mathematical details for how jumps, flows and switches are computed. For completeness, we redefine equations from the main text here. In particular, we recall that the hybrid JSF model we propose formally takes the following form, for any time interval $t_{0}<t<t_{1}$ between switching events,
$\frac{\mathrm{d}\vec{V}_{F}}{\mathrm{d}t}$ $=\eta_{\mathcal{S}^{\prime}}\vec{\lambda}_{\mathcal{S}^{\prime}}(\vec{V})+\bar{\eta}_{\mathcal{S}}\vec{\Lambda}_{\mathcal{S}}(\vec{V}),$ (S1)
$\vec{V}_{J}(t)$ $=\vec{V}_{J}(t_{0})+\eta_{\mathcal{S}}\int_{t_{0}}^{t}\vec{\Lambda}_{\mathcal{S}}(\vec{V})\ \mathrm{d}s,$ (S2)
where $\vec{\lambda}_{\mathcal{S}^{\prime}}\in\mathbb{R}^{p}$ are the reaction rates of flows and $\vec{\Lambda}_{\mathcal{S}}$ is a stochastic vector of $m-p$ delta-function spike trains that are derived from the realisations of $m-p$ different jumps sampled at rates which are dependent on the dynamic changes in the propensities $\vec{\lambda}_{\mathcal{S}}\in\mathbb{R}^{m-p}$ for these jumps. We recall that $\eta\in\mathbb{Z}^{n,m}$ is the stoichiometric matrix, written as:
$\eta=\left(\begin{array}[]{c|c}\eta_{\mathcal{S}^{\prime}}&\bar{\eta}_{\mathcal{S}}\\ \hline\cr 0&\eta_{\mathcal{S}}\end{array}\right),$ (S5)
where $\eta_{\mathcal{S}^{\prime}}\in\mathbb{Z}^{q\times p}$, $\eta_{\mathcal{S}}\in\mathbb{Z}^{(n-q)\times(m-p)}$ and $\bar{\eta}_{\mathcal{S}}\in\mathbb{Z}^{q\times(m-p)}$ refer to stoichiometric coefficients for changes in flowing compartments under flows, jumping compartments under jumps, flowing compartments under jumps, respectively, and $q\leq n$ is the number of flowing compartments.
S1.1 Jump events
The reactions that are defined in $\mathcal{S}$ are stochastic events that produce discontinuous jumps in the state vector $\vec{V}$. In (S1)-(S2), we denote the jumps using the notation $\vec{\Lambda}_{\mathcal{S}}(\vec{V})$. Each element in the vector $\vec{\Lambda}_{\mathcal{S}}$ corresponds to a reaction in $\mathcal{S}$. Consider, for example, $\mathcal{S}_{K}$ the $K$th reaction in $\mathcal{S}$ and we shall suppose that this reaction corresponds to the $k$th reaction in the full model $\mathcal{S}_{K}=\mathcal{R}_{k}$. The $K$th element of $\vec{\Lambda}_{\mathcal{S}}$ is
$\Lambda_{\mathcal{S},K}=\sum_{e}\delta(t-t_{k}^{(e)}),$
where $\delta$ is the Dirac measure and $t_{k}^{(e)}$ is the $e$th time at which the jump event $\mathcal{R}_{k}$ takes place. In this way, the term $\vec{\Lambda}_{\mathcal{S}}$ in (S1)-(S2) manifests as discrete jumps in both $\vec{V}_{F}$ and $\vec{V}_{J}$ at the moments of each jump event. For the sake of the simulation, the computation of the jump times $t_{k}^{(e)}$ for each instance $e$ of each reaction $\mathcal{R}_{k}\in\mathcal{S}$ is all that is required. The stoichiometric coefficients present in (S1)-(S2) then indicate the amplitude of the jumps in each compartment.
[…]
$J_{k}(t_{k}):=\log(u_{k}^{-1})-\int_{t_{0}}^{t_{k}}\lambda_{k}(\vec{V}(s))\ \mathrm{d}s=0.$ (S6)
In general, we cannot solve directly for $t_{k}$. Instead, we solve for it numerically by tracking the value of $J_{k}(t)$ as $\vec{V}$ evolves through flows, jumps and switches. For each reaction $\mathcal{R}_{k}$, at some initial time (for example $t_{k}^{(e-1)}$) we sample $u_{k}$ and initialise $t_{0}=t_{k}^{(e-1)}$. The initial value of $J_{k}(t)$ is therefore equal to the positive number $\log(u_{k}^{-1})$. As time progresses, $J_{k}(t)$ decreases according to (S6) since $\lambda_{k}\geq 0$. Its value ‘ticks’ down to zero over time and when $J_{k}$ reaches 0, a jump associated with $\mathcal{R}_{k}$ is triggered (hence the name jump clock). As a jump clock reaches 0 and a jump is triggered, the clock is reset by sampling a new random number ,$u_{k}\sim\text{Unif}(0,1)$. To update the jump clock, we require numerical integration of $\lambda_{k}(\vec{V}(t))$ forward in time. Fortunately, we also have piece-wise polynomial approximations for $\vec{V}_{F}(t)$ as a result of our numerical treatment of the continuous flows (see Subsection S1.2 ‘Flow events’) combined with piece-wise constant values for $\vec{V}_{J}(t)$ which only update once jumps occur. We will discuss further the numerical integration and jump clock updates in Subsection S1.3 ‘Jump clock updates’.
S1.2 Flow events
Between jumps, $\vec{V}_{J}$ remains constant and $\vec{V}_{F}$ evolves continuously according to (S1). In particular,
$\frac{\mathrm{d}\vec{V}_{F}}{\mathrm{d}t}=\eta_{\mathcal{S}^{\prime}}\vec{\lambda}_{\mathcal{S}^{\prime}}(\vec{V}).$ (S7)
This is a standard dynamical system of ODEs. We shall numerically integrate (S7) forward in time over discrete time steps $\Delta t$ using a simple Forward Euler method. However, higher order forward methods may be substituted. In particular,
$\vec{V}_{F}(t+\Delta t)=\vec{V}_{F}(t)+\Delta t\,\Delta\vec{V}_{F}(t)=\vec{V}_{F}(t)+\Delta t\,\eta_{\mathcal{S}^{\prime}}\vec{\lambda}_{\mathcal{S}^{\prime}}(\vec{V}(t)),$ (S8)
noting that $\vec{V}_{J}(t+\Delta t)=\vec{V}_{J}(t)$.
S1.3 Jump clock updates
We return to (S6). For a given jump reaction $\mathcal{R}_{k}$, a clock is initialised at time $t_{0}$; and $u_{k}\sim\text{Unif}(0,1)$, giving $J_{k}=\log(u_{k}^{-1})$. Over the course of a time step from $t$ to $t+\Delta t$, it is observed from (S6) that the clock ticks down from $J_{k}$ to $J_{k}-\Delta J_{k}$ where, since $\Delta t$ is small
$\Delta J_{k}=\int_{0}^{\Delta t}\lambda_{k}(\vec{V}(t+s))\ \mathrm{d}s$ $=\int_{0}^{\Delta t}\left(\alpha+\beta s+O(s^{2})\right)\ \mathrm{d}s,$ (S9)
$\approx\frac{\Delta t}{2}\left(2\alpha+\beta\Delta t\right),$ (S10)
$=\frac{\Delta t}{2}\left(2\alpha+(\Delta t\Delta\vec{V}^{\intercal}_{F})(\nabla_{\vec{V}_{F}}\lambda_{k})\right),$ (S11)
where $\alpha=\lambda_{k}(\vec{V}(t))$ is simply the propensity of reaction $\mathcal{R}_{k}$ at time $t$ and $\beta=\dfrac{\mathrm{d}\vec{V}^{\intercal}(t)}{\mathrm{d}t}\nabla_{\vec{V}}\lambda_{k}$. Importantly, we know that between jumps $\vec{V}^{\prime}_{F}$ is given by (S7) whilst $\vec{V}^{\prime}_{J}=0$. Thus, $\beta=\vec{\lambda}^{\intercal}_{\mathcal{S}^{\prime}}(\vec{V})\eta^{\intercal}_{\mathcal{S}^{\prime}}\nabla_{\vec{V}_{F}}\lambda_{k}$ (evaluated at $t$). Equation (S11) then follows from (S8).
To calculate the updated jump clock, we subtract $\Delta J_{k}$ from $J_{k}(t)$ to get a provisional $J_{k}(t+\Delta t)$. If $J_{k}(t)-\Delta J_{k}>0$, then no jump occurred during the interval $(t,t+\Delta t)$ and we have the jump clock $J_{k}(t+\Delta t):=J_{k}(t)-\Delta J_{k}$. If instead $J_{k}(t)-\Delta J_{k}<0$ then there is a jump (i.e. a $\mathcal{R}_{k}$ reaction) during the interval $(t,t+\Delta t)$ (where $0<\Delta\tau<\Delta t$) which we need to account for in the updated jump clock. Let $t+\Delta\tau$ denote the time at which this jump occurs. We can find $\Delta\tau$ by solving the equation $02\Delta J_{k}-\Delta\tau(2\alpha+\beta\Delta\tau)=$ 0 where $\Delta J_{k}$ is the residual of the jump clock from $t$ to $t+\Delta\tau$. However, the time $t+\Delta\tau$ ($0<\Delta\tau<\Delta t$) where the jump occurs can be found by solving $02\Delta J_{k}-\Delta\tau(2\alpha+\beta\Delta\tau)=$ 0 where $\Delta J_{k}$ is the residual of the jump clock from $t$ to $t+\Delta\tau$.
$\Delta\tau=\begin{cases}\frac{\sqrt{\alpha^{2}+2\beta\Delta J_{k}}-\alpha}{\beta},\quad&\beta\neq 0,\\ \frac{\alpha}{\Delta J_{k}},\quad&\beta=0.\end{cases}$ (S12)
In the case that a jump clock runs out, instead of using $\Delta t$ to push forward the flows in Equation (S8) we instead use $\Delta\tau$ and implement the jump after the flow to time $t+\Delta\tau$. Subsequently, we reinitialise the jump clock $J_{k}$. For a summary of this procedure see Figure S1.
Figure S1: The jump clock $J_{k}$ counts down until there is a jump change in the state. The jump event occurs between $t$ and $t+\Delta t$. To implement, we solve for the time $t+\Delta\tau$ using (S12), we then adjust the flows by using $\Delta\tau$ as a time step instead of $\Delta t$ in (S8). Once the jump has occurred, we reinitialise $J_{k}=\log(u_{k}^{-1})$ where $u_{k}\sim\text{Unif}(0,1)$.
S1.4 Switching events
Switching events describe instances where compartment membership of $\vec{V}_{J}$ and $\vec{V}_{F}$ can suddenly change as well as reaction membership in $\mathcal{S}$ and $\mathcal{S}^{\prime}$. There are two types of switching events. The first involves the transitioning of a compartment from $\vec{V}_{J}$ to $\vec{V}_{F}$. This transition is straightforward as a new equation is added to (S7) the state $\vec{V}_{F}$ is initialised at the switching time by continuation and initialisation of the new flowing compartment at $\Omega$.
The second type of switching event involves the transition of a compartment from $\vec{V}_{F}$ to $\vec{V}_{J}$. Let $V_{i}$ be the compartment switching from $\vec{V}_{F}$ to $\vec{V}_{J}$ due to a jump event, such that $V_{i}\leq\Omega$. In general, these types of jump events result in $V_{i}$ being non-integer, i.e. $V_{i}\notin\mathcal{V}_{\Omega}$. To ensure the values of $V_{i}$ stay in $\mathcal{V}_{\Omega}$ we add another constraint to the process proposed by (rebuli2017hybrid). Let $\hat{V}_{i}$ be the value of the flowing compartment $V_{i}$ after jumping down across the threshold $\Omega$ but before being initialised into $\mathcal{V}_{\Omega}$. We apply the following rule to reinitialise $V_{i}$ after the switch. We take $V_{i}=\lceil\hat{V}_{i}\rceil$ with probability $\hat{V}_{i}-\lfloor\hat{V}_{i}\rfloor$, and otherwise we round down by setting $V_{i}=\lfloor\hat{V}_{i}\rfloor$. This ensures the expected state of the process after the switch is $\hat{V}_{i}$ as described under the flowing paradigm from which this compartment has come and that the variable remains in the domain $\mathcal{V}_{\Omega}$.
S2 Jump-Switch-Flow: implementation details
@@
S2.1 Jump-clock
We implement the JSF algorithm with a slightly different jump clock as defined by $J_{k}$ in Section S1.1. In particular, equate $u_{k}=\text{CDF}(t_{k};k)$ but do not take the log of this expression before defining $J_{k}$ in Equation (S6). As a result, in our codes we consider the jump clock $\tilde{J}_{k}$:
$\tilde{J}_{k}(t)=u_{k}-\left(1-\exp\left\{-\int_{t_{0}}^{t}\lambda_{k}(\vec{V}(s))\ \mathrm{d}s\right\}\right),$ (S13)
where $u_{k}\sim\text{Unif}(0,1)$ and the time of the previous jump event $t_{0}$. Computing the times for the jump events follows a similar process as above. We compute the jump clock on the regular ODE mesh used to solve $\vec{V}_{F}$ in Equation (S7). From Equation (S13) we can write $\tilde{J}_{k}(t+\Delta t)$ in terms of $\tilde{J}_{k}(t)$ and an integral of the reaction rate:
$\tilde{J}_{k}(t+\Delta t)=\tilde{J}_{k}(t)+\left(\exp\left\{-\int_{t}^{t+\Delta t}\lambda_{k}(\vec{V}(s))\ \mathrm{d}s\right\}-1\right)\left(\tilde{J}_{k}(t)+1-u_{k}\right).$ (S14)
For steps where $\tilde{J}_{k}(t+\Delta t)>0$ we can do a single step of this process to get the updated jump clock values. However, if $\tilde{J}_{k}(t+\Delta t)<0$ a jump has occurred at some time $t+\Delta\tau\in\left(t,t+\Delta t\right)$, and we need to solve $\tilde{J}_{k}(t+\Delta\tau)=0$ to find when the jump occurred. Therefore, we require a method to find $\Delta\tau$, where $0<\Delta\tau<\Delta t$. To do this, we start by letting the right hand side of Equation (S14) equal zero and rearrange to get:
$\int_{t}^{t+\Delta\tau}\lambda_{k}(\vec{V}(s))\ \mathrm{d}s=\ln\left\{\frac{\tilde{J}_{k}(t)+1-u_{k}}{1-u_{k}}\right\}.$ (S15)
@@
$\Delta\tau=\frac{1}{\lambda_{k}}\ln\left\{\frac{\tilde{J}_{k}(t)+1-u_{k}}{1-u_{k}}\right\}.$ (S16)
If $\lambda_{k}$ is not constant, we approximate $\lambda_{k}$, which we call $\widehat{\lambda}_{k}$, using the integration of $\vec{V}(t)$, as per Equation (S7), since we know $\vec{V}(t)$ and $\vec{V}(t+\Delta t)$. Therefore, we also know $\lambda_{k}$ at $t$ and $t+\Delta t$. We then write $\widehat{\lambda}_{k}$, at $t+\Delta\tau$ between times times $t$ and $t+\Delta t$ via linear interpolation:
$\widehat{\lambda}_{k}(t+\Delta\tau)\approx\left(1-\frac{\Delta\tau}{\Delta t}\right)\lambda_{k}(\vec{V}(t))+\frac{\Delta\tau}{\Delta t}\lambda_{k}(\vec{V}(t+\Delta t)).$ (S17)
We can now substitute this approximation into Equation (S15), solve it, and rearrange for $\Delta\tau$. This gives a quadratic in $\Delta\tau$ with the following solution:
$\Delta\tau=\frac{\sqrt{\left(\Delta t\widehat{\lambda}_{k}(t)\right)^{2}+2\alpha\Delta t\Delta\lambda_{k}}-\Delta t\widehat{\lambda}_{k}(t)}{\Delta\lambda_{k}},$ (S18)
@@
Here, we describe an exact algorithm for sampling trajectories from Jump-Switch-Flow. The algorithm to sample exactly from Jump-Switch-Flow is given in Algorithm S1. First, initialise the jump clocks (line 3). On each iteration through the ODE loop, we first partition the reactions based on inclusion in $\mathcal{S}$ (line 6). We use the IVP solver to compute $\Delta\vec{V}_{F}(t)$ based on the flowing reactions (line 7). For each step of the ODE loop, we update the jump clocks (line 9). On lines 10–12, we determine if a jump event has occurred based on the sign of the the Jump-clock. If a jump event has occurred, compute the corresponding event time. On line 13 we then rewind the state (and jump clocks) to when the jump occurred. Resample the appropriate jump clock (line 14), and update the state to account for the jump (line 15). Repeat until the time has reached the maximum desired time.
Algorithm S1 An algorithmic description of the Exact Jump-Switch-Flow sampler.
0:  $\vec{V}(0)$, $\Delta t>0$ and $T_{\max}>0$

### v3

@@ Results
Recall that the particle filter is capable of combining a mechanistic model of the within-host dynamics, such as the TEIRV, and an observation model, such as the viral load measurements, and will return a (Bayesian) posterior sample of both the parameters of the process and the trajectory of virus and cell populations through time. We assume that the infection begins with a single exposed target cell (in a population of $8\times 10^{7}$ target cells) [23]. This gives us an initial condition for the process: $T(0)=8\times 10^{7}$, $E(0)=1$, $I(0)=0$, $R(0)=0$. We leave the initial viral load, $V(0)$, as a parameter to be fit to the data.
We selected six patient time series [23], choosing ones that contained a full 14 data points, for both consistency and simplicity. Two of the model parameters were fixed as in previous analysis [23]: $c=10$, $k=4$. See SI Section 3 for full details of the particle filter and JSF configuration.
@@
Figure 5: JSF enables the inference of the viral load through time and the probability that the virus has been cleared from longitudinal data of SARS-CoV-2 viral load measurements. A Model fits to viral load data based on nasal swabs of six patients from [23] using a refractory cell within-host model, using our Jump-Switch-Flow method. Nasal viral load data is shown as solid dots, and the limit of detection is shown as a dotted line. We show the switching threshold, $\Omega=10^{2}$ by a dashed line. The red and purple bands show mean (0%), 25%, 50%, 75% and 95% credible intervals (from dark to light) for the inference and prediction, respectively. We also estimate the probability of viral clearance at a given point in time, based on the previous data to that point, in blue (light). B The posterior distribution of viral $R_{0}$ for each patient, along with the median (white circle), These patient specific estimates are contrasted with the single point estimate from previous analysis [23].
We estimate the probability of viral clearance over time (right blue axis of in the panels of Figure 5A). Viral clearance is defined by the point at which the virion, eclipsed, and infected cell populations all reach extinction. We note that this is a conservative measure if one were only interested in when a patient’s infectiousness becomes negligible. However, for proof of concept of inferring such quantities, we have chosen to estimate the probability of complete viral clearance.
There is clear heterogeneity in the amount of time required to clear the virus (Figure 5A). We estimate that with high certainty patients 423192, 443108 and 445602 clear the virus within 20 days. In contrast, for patients 444332, 444391 and 451152 we infer that they have not cleared the virus within that time frame. Within those who did clear the virus, we observe that different patients require different amounts of time to clear the virus. Specifically, patient 432192 (5A) is inferred to have cleared the virus first. Patients 443108 and 445602 both require an estimated 16 days to obtain viral clearance.
The posterior distributions of the parameters (Figure SI 11) and priors (Table SI 5), are given in SI Section 5. Because of potential identifiability issues with the rate parameters, we instead compare the estimated reproduction number, $R_{0}$, to results from previous analysis [23]. Figure 5B shows the posterior distributions of $R_{0}$, as well as the point estimates from [23]. Since the previous analysis, [23], reports identical estimates for $\pi$, $\beta$ and $\delta$ across these patients, we only include a single point estimate for their work. Our estimate is consistent with previous results [23] for patients 432192, 445602 and 451152. However, patients 443108, 444332 and 444391 all have higher $R_{0}$ estimates in our results. Our $R_{0}$ estimates are consistent with similar within-host viral infection analyses of other respiratory pathogens [5, 20, 19].
@@ Discussion
We have presented a simple hybrid simulation method for compartmental models: Jump-Switch-Flow (JSF). This method facilitates efficient simulation of multi-scale models. Via simulation studies and an analysis of longitudinal data from SARS-CoV-2 infections we demonstrated the desirable computational properties of this method and the types of novel analyses which it enables. JSF allows compartments to dynamically change between stochastic and deterministic behaviour. Combining JSF with a simulation based inference method, e.g. a particle filter, enables inference for multi-scale models, and in situations where absorbing states (such as a population going extinct) are important.
@@
Simulations shows JSF produces trajectories largely consistent with gold standard exact simulation techniques, e.g. Doob-Gillespie algorithm (see for example Figures SI 1–SI 4 and SI 6 – SI 9). We demonstrate how JSF is highly accurate at producing summary statistics comparable to gold standard exact simulation techniques (see Figures SI 1 – SI 4), with a significant improvement in accuracy for problems with multiple time scales when compared to currently available hybrid methods (see Figure SI 3 and SI 4) However, JSF is also much faster for realistic population sizes (as can be seen in the comparison of computational speed in Figure SI 10B). We demonstrate how our approach can be incorporated into a particle filter [35], to perform parameter and state estimation and prediction (see Figure 3).
A strength of JSF is the capacity to infer when a process has reached an absorbing state, for example, an epidemic fading-out, or a virus being cleared. The ability to infer when an absorbing state has been reached is important for calibrating intervention measures [36] and understanding immune response [44]. ODE based models rarely have absorbing states that can be reached in finite time, which fundamentally limits their capacity to describe extinction and clearance processes.
@@
We analysed viral load for SARS-CoV-2 infections using the JSF within a particle filter with a TEIRV, refractory cell model [23]. In the subset of the data considered, we find consistent parameter values between patients (SI Section 3). In a novel analysis, we estimated the probability of viral clearance through time, finding substantial heterogeneity in the time until viral clearance (see Figure 5A). This is important as an accurate quantification of when the virus has been cleared is crucial for determining appropriate treatment regimes.
@@
A key advantage of JSF is the ability to combine the stochasticity of discrete small populations with convenient deterministic models for large populations. The point at which a compartment will transition between these descriptions is specified by the threshold parameters $\Omega_{i}$ (one for each compartment). Setting these parameters to low values speeds up computation, while setting them higher captures more of the stochasticity in the process. As demonstrated in the simulation studies, an appropriate value can be determined through some preliminary simulations (see SI 4 for an example of preliminary simulation analysis for the SIRS model example). An important consideration in selecting this parameter is ensuring that the absorbing states can still be reached (e.g. compartment extinction), and that stable and steady states can be perturbed by random fluctuations. Potential considerations when specifying the values $\Omega_{i}$ may also include preliminary simulations, increasing $\Omega_{i}$ and observing how the computational complexity scales (see Figure SI 10). Furthermore, developing theory to better understand when a CTMC is well approximated by an ODE is at the core to determine how the switching threshold should be chosen. Understanding into the numerical properties of the Jump-Switch-Flow process, and how convergence of the method behaves for arbitrary choices of Flow event solvers will also prove to be an exciting area of further research. However, these lies outside the scope of this work.
Extensions include the incorporation of an intermittent Stochastic Differential Equation approximation between the CTMC and ODE regimes, and extension to spatial processes which exhibit both stochastic and deterministic behaviours.
@@ Acknowledgement
We thank Ada Yan for helpful discussions regarding within-host modelling. J.A.F.’s research is supported by the Australian Research Council (FT210100034, CE230100001) and the National Health and Medical Research Council (APP2019093). S.H.’s research is supported by the Australian Research Council (DP200101281).
Data Availability
The manuscript has associated data in a data repository available at https://github.com/DGermano8/JSFGermano2024. Original data related to SARS-CoV-2 viral load measurements can also be found at [23].
Code Availability
All code to reproduce our results is available at https://github.com/DGermano8/JSFGermano2024. A Python package implementing the Jump-Switch-Flow method is available at https://DGermano8.github.io/JSF.
@@
J.A.F and M.B.F conceptualised research; D.P.J.G, A.E.Z, S.H., R.M, J.A.F and M.B.F designed research; D.P.J.G and A.E.Z performed research; D.P.J.G, A.E.Z and R.M contributed software tools; D.P.J.G and A.E.Z analyzed data; J.A.F and M.B.F supervised research; D.P.J.G and A.E.Z wrote the paper; D.P.J.G, A.E.Z, S.H., R.M, J.A.F and M.B.F reviewed the paper; J.A.F acquired funding.
Supplementary Information: A hybrid framework for compartmental models enabling simulation-based inference
@@
SI1 Jump-Switch-Flow: implementation details
@@
SI1.1 Jump-clock
We implement the JSF algorithm with a slightly different jump clock as defined by $J_{k}$ in Section Jump clock updates. In particular, equate $u_{k}=\text{CDF}(t_{k};k)$ but do not take the log of this expression before defining $J_{k}$ in Equation (13). As a result, in our codes we consider the jump clock $\tilde{J}_{k}$:
$\tilde{J}_{k}(t)=u_{k}-\left(1-\exp\left\{-\int_{t_{0}}^{t}\lambda_{k}(\vec{V}(s))\ \mathrm{d}s\right\}\right),$ (SI1)
where $u_{k}\sim\text{Unif}(0,1)$ and the time of the previous jump event $t_{0}$. Computing the times for the jump events follows a similar process as above. We compute the jump clock on the regular ODE mesh used to solve $\vec{V}_{F}$ in Equation (8). From Equation (SI1) we can write $\tilde{J}_{k}(t+\Delta t)$ in terms of $\tilde{J}_{k}(t)$ and an integral of the reaction rate:
$\tilde{J}_{k}(t+\Delta t)=\tilde{J}_{k}(t)+\left(\exp\left\{-\int_{t}^{t+\Delta t}\lambda_{k}(\vec{V}(s))\ \mathrm{d}s\right\}-1\right)\left(\tilde{J}_{k}(t)+1-u_{k}\right).$ (SI2)
For steps where $\tilde{J}_{k}(t+\Delta t)>0$ we can do a single step of this process to get the updated jump clock values. However, if $\tilde{J}_{k}(t+\Delta t)<0$ a jump has occurred at some time $t+\Delta\tau\in\left(t,t+\Delta t\right)$, and we need to solve $\tilde{J}_{k}(t+\Delta\tau)=0$ to find when the jump occurred. Therefore, we require a method to find $\Delta\tau$, where $0<\Delta\tau<\Delta t$. To do this, we start by letting the right hand side of Equation (SI2) equal zero and rearrange to get:
$\int_{t}^{t+\Delta\tau}\lambda_{k}(\vec{V}(s))\ \mathrm{d}s=\ln\left\{\frac{\tilde{J}_{k}(t)+1-u_{k}}{1-u_{k}}\right\}.$ (SI3)
@@
$\Delta\tau=\frac{1}{\lambda_{k}}\ln\left\{\frac{\tilde{J}_{k}(t)+1-u_{k}}{1-u_{k}}\right\}.$ (SI4)
If $\lambda_{k}$ is not constant, we approximate $\lambda_{k}$, which we call $\widehat{\lambda}_{k}$, using the integration of $\vec{V}(t)$, as per Equation (8), since we know $\vec{V}(t)$ and $\vec{V}(t+\Delta t)$. Therefore, we also know $\lambda_{k}$ at $t$ and $t+\Delta t$. We then write $\widehat{\lambda}_{k}$, at $t+\Delta\tau$ between times times $t$ and $t+\Delta t$ via linear interpolation:
$\widehat{\lambda}_{k}(t+\Delta\tau)\approx\left(1-\frac{\Delta\tau}{\Delta t}\right)\lambda_{k}(\vec{V}(t))+\frac{\Delta\tau}{\Delta t}\lambda_{k}(\vec{V}(t+\Delta t)).$ (SI5)
We can now substitute this approximation into Equation (SI3), solve it, and rearrange for $\Delta\tau$. This gives a quadratic in $\Delta\tau$ with the following solution:
$\Delta\tau=\frac{\sqrt{\left(\Delta t\widehat{\lambda}_{k}(t)\right)^{2}+2\alpha\Delta t\Delta\lambda_{k}}-\Delta t\widehat{\lambda}_{k}(t)}{\Delta\lambda_{k}},$ (SI6)
@@
Here, we describe an exact algorithm for sampling trajectories from Jump-Switch-Flow. The algorithm to sample exactly from Jump-Switch-Flow is given in Algorithm SI1. First, initialise the jump clocks (line 3). On each iteration through the ODE loop, we first partition the reactions based on inclusion in $\mathcal{S}$ (line 6). We use the IVP solver to compute $\Delta\vec{V}_{F}(t)$ based on the flowing reactions (line 7). For each step of the ODE loop, we update the jump clocks (line 9). On lines 10–12, we determine if a jump event has occurred based on the sign of the the Jump-clock. If a jump event has occurred, compute the corresponding event time. On line 13 we then rewind the state (and jump clocks) to when the jump occurred. Resample the appropriate jump clock (line 14), and update the state to account for the jump (line 15). Repeat until the time has reached the maximum desired time.
Algorithm SI1 An algorithmic description of the Exact Jump-Switch-Flow sampler.
$\vec{V}(0)$, $\Delta t>0$ and $T_{\max}>0$